# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

### v1

@@ 3 Asymptotics for estimators: a theoretical foundation
Let $A\subset{\mathbb{R}}^{d}$ be a nonempty closed set. The topological boundary of $A$ is denoted by $\partial A$. The (positive) normal bundle of $A$ is the subset of ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$ which is defined by
$\operatorname{Nor}(A):=\{(x,u)\in\partial A\times\mathbb{S}^{d-1}:{\rm dist}(A,x+ru)=r\text{ for some }r>0\}$
and the reach function of $A$ is given by
@@
(note the slight correction in comparison with [23], see also [24]). In the following, we write $\mathcal{H}^{k}$ for the $k$-dimensional Hausdorff measure for $k\in\{0,\ldots,d\}$. We remark that $A$ has reach at least $r>0$ if $r_{A}(x,u)\geq r$ for $\mathcal{H}^{d-1}$-almost all $(x,u)\in\operatorname{Nor}(A)$; see [24, Lemma 3.19] for a more general fact.
It was shown in [23, Theorem 2.1], [24, Theorem 3.16] that there exist uniquely determined signed “reach measures” $\mu_{i}(A;\cdot)$ on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$, for $i\in\{0,\ldots,d-1\}$, concentrated on $\operatorname{Nor}(A)$ and such that
@@
where $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is an arbitrary measurable bounded function with compact support. Clearly, (3) remains true if $f$ takes values in the (finite-dimensional) vector space $\mathbb{T}^{r}({\mathbb{R}}^{d})$, for any fixed $r\in{\mathbb{N}}_{0}$, where $f:{\mathbb{R}}^{d}\to\mathbb{T}^{r}({\mathbb{R}}^{d})$ is said to be bounded if the map $x\mapsto|f(x)|$ is bounded ($f$ is bounded on a subset $B$ if the restriction of this map to $B$ is bounded).
@@
If $A\in\mathcal{U}^{d}$, then the support measures $\Lambda_{j}(A,\cdot)$ of $A$ are defined by additive extension. For this generalization it is crucial that $A$ can be obtained as a finite union of compact sets $A_{i}$ with positive reach, $i\in\{1,\ldots,m\}$, such that arbitrary finite intersections of these sets have again positive reach (compare [49, Example 4.16]). It was shown in [23, Section 3] that
@@
Moreover, the arguments in [23, Sections 3 and 4] show that indeed
@@
as Borel measures on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$; see also [48, Corollary 2]. In particular, $A\mapsto\mu_{d-1}(A;\cdot)$ is additive on $\mathcal{U}^{d}$. On the other hand, the support measures are in general not defined on the larger domain of finite unions of sets with positive reach. For this reason, we work with the reach measures whenever the support measures are not available. Relation (3.2) shows that this approach is consistent with the previous literature if we restrict ourselves to the top order measures with index $j=d-1$ as in definition (4.2) below.
The integrals on the right-hand side of (3) are well defined, but in general the reach measures $\mu_{i}(A;\cdot)$ are not Radon measures. In the following, we consider compact sets $A\subset{\mathbb{R}}^{d}$ and assume that the (nonnegative) total variation measure $\|\mu_{i}(A)\|(\cdot)$ of $\mu_{i}(A;\cdot)$ is finite for $i\in\{0,\ldots,d-1\}$. A sufficient condition is provided in Lemma 3.2.
@@
If $A$ is a finite union of compact sets of positive reach, then the total variation measures of the reach measures of $A$ are concentrated on $\operatorname{Nor}(A)$ and satisfy
@@
It follows from [23, Corollary 2.5] that
@@
where $c(d,i)$ is a finite constant and $|H_{d-1-i}(A,x,u)|\leq\binom{d-1}{i}$, as can be seen from [23, (2.13)]. Let $A=\bigcup_{i=1}^{m}A_{i}$ with compact sets $A_{i}$ with positive reach. It is well known that $\operatorname{Nor}(A_{i})$ is bilipschitz homeomorphic to the compact submanifold ${\rm dist}(A_{i},\cdot)^{-1}(\{r_{0}\})$, if $r_{0}$ is smaller than the reach of $A_{i}$. Hence $\mathcal{H}^{d-1}(\operatorname{Nor}(A_{i}))<\infty$. Since
$\operatorname{Nor}(A)\subset\bigcup_{i=1}^{m}\operatorname{Nor}(A_{i}),$
the assertion follows. ∎
Let $A\subset{\mathbb{R}}^{d}$ be a nonempty closed set. For $\mathcal{H}^{d}$-almost all $z\in{\mathbb{R}}^{d}\setminus A$, the metric projection $p_{A}(z)\in A$ to $A$ and the vector $u_{A}(z)\in\mathbb{S}^{d-1}$ are uniquely determined by
$\|p_{A}(z)-z\|={\rm dist}(A,z):=\min\{\|a-z\|:a\in A\},\quad u_{A}(z)=\frac{z-p_{A}(z)}{{\rm dist}(A,z)}.$
Note that $p_{A}(z)=z$ for $z\in A$, and if $(x,u)\in\operatorname{Nor}(A)$ and $0<s<r_{A}(x,u)$, then $p_{A}(x+su)=x$ and $u_{A}(x+su)=u$. If $A$ is a nonempty compact set, we consider
@@
Note that $N(A,x)$ is either a singleton, consists of two unit vectors {±$\{\pm u\}$ or is an (infinite) spherically convex set. Then we set
$\partial^{i}A:=\{x\in\partial^{+}A:|N(A,x)|=i\},\quad i\in\{1,2\}.$
We write $\nu_{A}(x)$ for the unique unit vector in $N(A,x)$, for $x\in\partial^{1}A$, and ±$\pm\nu_{A}(x)$ for the two unit vectors in $N(A,x)$, if $x\in\partial^{2}A$. For $\varepsilon>0$, let $A^{\varepsilon}:=\{x\in{\mathbb{R}}^{d}:{\rm dist}(A,x)\leq\varepsilon\}$ denote the $\varepsilon$ neighborhood of $A$.
@@
Let $r,s\in\mathbb{N}_{0}$. Let $A\subset{\mathbb{R}}^{d}$ be a nonemtpy compact set such that (3.3) holds. If $f:{\mathbb{R}}^{d}\to\mathbb{T}^{r}({\mathbb{R}}^{d})$ is measurable and bounded on $\partial A$, then
@@
We apply (3) with the map $\tilde{f}:{\mathbb{R}}^{d}\to\mathbb{T}^{r+s}$, $\tilde{f}(z)=f(p_{A}(z))u_{A}(z)^{s}\mathds{1}\{z\in A^{\varepsilon}\setminus A\}$, which yields
@@
$\leq\frac{\omega_{d-i}}{d-i}\cdot|f|_{\partial A}\cdot\|\mu_{i}(A)\|(\operatorname{Nor}(A))\cdot\varepsilon^{d-i-1}\to 0\quad\text{ as }\varepsilon\to 0_{+},$
where $|f|_{\partial A}:=\sup\{|f(x)|:x\in\partial A\}<\infty$ and basic properties of the norm ||$\cdot$|| ⋅ | on (symmetric) tensors were used, in particular (2.2). Moreover, by the dominated convergence theorem,
@@
since $r_{A}(x,u)>0$ for $(x,u)\in\operatorname{Nor}(A)$. Note that by [23, Proposition 4.1], $\mu_{d-1}(A;\cdot)$ is nonnegative and
@@
By essentially the same argument, we obtain the following variant of Theorem 3.3. The case $s=0$ of the next theorem is already covered by Theorem 3.3 with the domain of integration restricted to the complement of the set $A$.

### v2

@@ 3 Asymptotics for estimators: a theoretical foundation
In the current section, we first recall and illustrate a local Steiner formula for general compact sets together with the required concepts from geometric measure theory. For special classes of sets, we explain how the reach measures involved are related to the better known support measures. The main results of this section are Theorems 3.7 and 3.9. They relate differentials of local volume integrals to surface integrals for general classes of compact sets, including finite unions of sets with positive reach. These results provide the foundation for the Voronoi-FD algorithm introduced in Section 4.2.
Let $A\subset{\mathbb{R}}^{d}$ be a nonempty closed set. The topological boundary of $A$ is denoted by $\partial A$. For $z\in{\mathbb{R}}^{d}$, the distance of $z$ from $A$ is denoted by $\text{dist}(A,z):=\min\{\|a-z\|:a\in A\}$. The $r$-neighborhood of $A$, for $r\geq 0$, is the set
$A^{r}:=\{y\in{\mathbb{R}}^{d}:\text{dist}(A,y)\leq r\}.$
The (positive) normal bundle of $A$ is the Borel subset of ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$ that is defined by
$\operatorname{Nor}(A):=\{(x,u)\in\partial A\times\mathbb{S}^{d-1}:{\rm dist}(A,x+ru)=r\text{ for some }r>0\},$
and the (Borel measurable) reach function of $A$ is given by
@@
(see [29] and note the slight correction in comparison with [28]). For the preceding assertions regarding measurability, see the references in [29, Section 2.4].
Figure 3: Set AA with positive reach and various support elements $(x,u)\in\operatorname{Nor}(A)$.
The reach of $A$, denoted by $\operatorname{reach}(A)$, is defined as the supremum of all $r\geq 0$ such that each $z\in A^{r}$ has a unique nearest point in $A$. We say that $A$ has positive reach, if $\operatorname{reach}(A)>0$. It can be shown that $\operatorname{reach}(A)\geq r$, for some $r>0$, if $r_{A}(x,u)\geq r$ for $\mathcal{H}^{d-1}$-almost all $(x,u)\in\operatorname{Nor}(A)$; see [29, Lemma 3.19] for the proof of a more general fact. An illustration of elements of the normal bundle of a set with positive reach is given in Fig. 3, Fig. 4 illustrates the reach function for a polygonal arc which is a polyconvex set (but not a set with positive reach). Recall that $\mathcal{U}^{d}$ denotes the class of all finite unions of compact sets with positive reach all of whose finite intersections have again positive reach. The polygonal arc shown in Fig. 4 lies in $\mathcal{U}^{d}$.
It was shown in [28, Theorem 2.1], [29, Theorem 3.16] that for a general closed set $A\subset{\mathbb{R}}^{d}$ there exist uniquely determined signed “reach measures” $\mu_{i}(A;\cdot)$ on the Borel subsets of ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$, for $i\in\{0,\ldots,d-1\}$, concentrated on $\operatorname{Nor}(A)$ and such that
@@
where $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is an arbitrary measurable bounded function with compact support. Clearly, (3) remains true if $f$ takes values in the (finite-dimensional) vector space $\mathbb{T}^{r}({\mathbb{R}}^{d})$, for any fixed $r\in{\mathbb{N}}_{0}$, where $f:{\mathbb{R}}^{d}\to\mathbb{T}^{r}({\mathbb{R}}^{d})$ is said to be bounded if the map $x\mapsto|f(x)|$ is bounded ($f$ is bounded on a subset $B$ if the restriction of this map to $B$ is bounded). It is shown in [28, (2.27)] that on the right side of (3) the order of integration can be interchanged. Hence, for $(x,u)\in\operatorname{Nor}(A)$ the indicator $\mathds{1}\{r_{A}(x,u)>t\}$ restricts the domain of integration for $t\geq 0$ so that the triples $(x,u,t)$ uniquely parametrize almost all points of ${\mathbb{R}}^{d}$ via $(x,u,t)\mapsto x+tu\in{\mathbb{R}}^{d}$.
@@
If $A\neq\emptyset$ is a compact convex set, then $r_{A}(x,u)=\infty$ for $(x,u)\in\operatorname{Nor}(A)$ and (3) boils down to the local Steiner formula for convex bodies. If $\text{reach}(A)>r$ and the function $f$ is supported in $A^{r}$, then the indicator function in (3) can be omitted, and again we get a local Steiner formula in the $r$-neighborhood $A^{r}$ of $A$.
Figure 4: Boundary $A=\partial P$ of a polytope $P$ and particular values of the reach function with touching balls from outside.
Remark 3.2.
If $A\in\mathcal{U}^{d}$, then the support measures $\Lambda_{j}(A,\cdot)$, $j\in\{0,\ldots,d-1\}$, of $A$ are defined by additive extension. For this generalization, it is crucial that $A=A_{1}\cup\ldots\cup A_{m}$ can be obtained as a finite union of compact sets $A_{l}$ with positive reach, $l\in\{1,\ldots,m\}$, such that arbitrary finite intersections of $A_{1},\ldots,A_{m}$ have positive reach. In general, the intersection of two sets of positive reach does not have positive reach. A basic example is provided in [63, Example 4.16], an even more striking example is constructed in [5, Section 5, Example 1], another example is displayed in Fig. 5. The set $A$ is the epigraph of the function $t\mapsto f\left(\frac{\pi}{16}-|t|\right)$, $t\in[-\frac{\pi}{16},\frac{\pi}{16}]$, where $f(t):=t^{6}\cdot\sin^{2}\left(\frac{1}{t}\right)$, for $t\neq 0$, and $f(0):=0$.
Remark 3.3.
For a generic (in the sense of Baire categories [69, Section 2.6]) convex body $K\subset{\mathbb{R}}^{d}$ (even for $d=2$) the boundary $A=\partial K$ of $K$ is strictly convex, of class $C^{1}$, for $\mathcal{H}^{d-1}$-a.e. $x\in A$, $N(A,x):=\{u\in\mathbb{S}^{d-1}:(x,u)\in\operatorname{Nor}(A)\}$ consists of two antipodal unit vectors with $r_{A}(x,u)=\infty$ and $r_{A}(x,-u)\in(0,\infty)$ (say) and the set of points $x\in A$ where $r_{A}(x,-u)=0$ may be dense in $A$ (see [52] and Corollary 2.7.2, Theorem 2.7.4 and Note 2 for Section 2.7 in [69]). Clearly, $A$ is not a finite union of sets with positive reach.
Figure 5: The sets $A,B$ have positive reach, but neither $A\cap B$ nor $A\cup B$ has positive reach.
It was shown in [28, Section 3] that if $A\in\mathcal{U}^{d}$, then
@@
For $A\in\mathcal{U}^{d}$, the additive extension $\Lambda_{j}(A,\cdot)$ of the support measures is concentrated on $N^{*}(A)$, another notion of normal bundle that contains $\operatorname{Nor}(A)$ as a subset and is defined by means of an index function. The arguments in [28, Sections 3 and 4] show that indeed
@@
as Borel measures on ${\mathbb{R}}^{d}\times\mathbb{S}^{d-1}$; see also [62, Corollary 2]. In particular, $A\mapsto\mu_{d-1}(A;\cdot)$ is additive on $\mathcal{U}^{d}$. On the other hand, the support measures are in general not defined on the larger domain of finite unions of sets with positive reach. For this reason, we work with the reach measures whenever the support measures are not available. The relation (3.2) shows that this approach is consistent with the previous literature if we restrict ourselves to the top-order measures with index j=d1j=d-1, as in the definition (4.13) below.
The integrals on the right-hand side of (3) are well defined, but in general the reach measures $\mu_{i}(A;\cdot)$ are not Radon measures. In the following, we consider compact sets $A\subset{\mathbb{R}}^{d}$ and assume that the (nonnegative) total variation measure $\|\mu_{i}(A)\|(\cdot)$ of $\mu_{i}(A;\cdot)$ is finite for $i\in\{0,\ldots,d-1\}$. Sufficient conditions are provided in Lemma 3.4.
@@
If $A=\bigcup_{l=1}^{m}A_{l}$ with compact sets $A_{l}\subset{\mathbb{R}}^{d}$ such that $\mathcal{H}^{d-1}(\operatorname{Nor}(A_{l}))<\infty$, then the total variation measures of the reach measures of $A$ are concentrated on $\operatorname{Nor}(A)$ and satisfy
@@
In particular, (3.3) holds if $A$ is a finite union of compact sets of positive reach.
@@
It follows from [28, Corollary 2.5] that
@@
where $c(d,i)$ is a finite constant and $|H_{d-1-i}(A,x,u)|\leq\binom{d-1}{i}$, as can be seen from [28, (2.13)]. Let $A=\bigcup_{l=1}^{m}A_{l}$ with compact sets $A_{l}\subset{\mathbb{R}}^{d}$ such that $\mathcal{H}^{d-1}(\operatorname{Nor}(A_{l}))<\infty$. Since
$\operatorname{Nor}(A)\subset\bigcup_{l=1}^{m}\operatorname{Nor}(A_{l}),$
the second assertion follows.
If $A_{l}$ is a compact set with positive reach, then $\operatorname{Nor}(A_{l})$ is bilipschitz homeomorphic to the compact (d1)(d-1)-dimensional submanifold ${\rm dist}(A_{l},\cdot)^{-1}(\{r_{0}\})$ of ${\mathbb{R}}^{d}$, if $r_{0}<\operatorname{reach}(A_{l})$; see also Remark 4.4, Lemma 4.21 and Corollary 4.22 in [63]. Hence $\mathcal{H}^{d-1}(\operatorname{Nor}(A_{l}))<\infty$, which implies the remaining assertion. ∎
Remark 3.5.
If $A\neq\emptyset$ is compact and $\varepsilon>0$, then $A^{\varepsilon}$ satisfies condition (3.3); see [28, Corollary 4.4].
Let $A\subset{\mathbb{R}}^{d}$ be a nonempty closed set. For $\mathcal{H}^{d}$-almost all $z\in{\mathbb{R}}^{d}\setminus A$, the metric projection $p_{A}(z)\in A$ of $z$ to $A$ and the vector $u_{A}(z)\in\mathbb{S}^{d-1}$ are uniquely determined (see [28, (2.1)] and the references given there) by
$\|p_{A}(z)-z\|={\rm dist}(A,z),\quad u_{A}(z)=\frac{z-p_{A}(z)}{{\rm dist}(A,z)}.$
Note that $p_{A}(z)=z$ for $z\in A$ (whereas $u_{A}(z)$ is defined only for $z\notin A$), and if $(x,u)\in\operatorname{Nor}(A)$ and $0<s<r_{A}(x,u)$, then $p_{A}(x+su)=x$ and $u_{A}(x+su)=u$. For a nonempty compact set $A\subset{\mathbb{R}}^{d}$, we consider
@@
Note that the Borel set $N(A,x)$ is either a singleton, consists of two unit vectors $\{\pm u\}$ or is an infinite spherically convex set. We set
$\partial^{i}A:=\{x\in\partial^{+}A:|N(A,x)|=i\},\quad i\in\{1,2\},$
where $|N(A,x)|$ denotes the cardinality of the set $N(A,x)$. We write $\nu_{A}(x)$ for the unique unit vector in $N(A,x)$, if $x\in\partial^{1}A$, and $\pm\nu_{A}(x)$ for the two unit vectors in $N(A,x)$, if $x\in\partial^{2}A$.
Remark 3.6.
The boundary $A=\partial P=\partial A$ of a polytope $P\subset{\mathbb{R}}^{2}$ is not a set with positive reach, but a finite union of (convex) segments. Note that $\partial^{+}A=\partial A$, $\partial^{1}A$ is the set of vertices and $\partial^{2}A$ is $\partial A$ minus the set of vertices. While here $A$ is still in the convex ring, this is no longer the case for the boundary of a general (planar, generic) convex body $K$, where still $\partial^{2}K$ has full measure in $\partial K$ and $\partial^{1}K$ can be a dense subset of $\partial K$ (see Remark 3.3).
An example of a set with positive reach that is the closure of its interior (but not a topological manifold) is provided in [5, Section 5, Example 1] (see the set between $A$ and $B$ in Fig. 5 for a simple illustration and compare with Remark 3.3). Clearly, these examples show the relevance of the sets $\partial^{1}A$ and $\partial^{2}A$.
The subsets $\partial^{1}A$ and $\partial^{2}A$ of $\partial A$ are crucial for the description of the limit in the following theorem. If $s$ is odd, the integral over $\partial^{2}A$ does not contribute, since $1+(-1)^{s}=0$. If $s$ is even, the integral is multiplied by $2$, which corresponds to the two unit normal vectors $\pm\nu_{A}(x)$ at boundary points $x\in\partial^{2}A$.
@@
Let $r,s\in\mathbb{N}_{0}$. Let $A\subset{\mathbb{R}}^{d}$ be a nonempty compact set such that (3.3) holds. If $f:{\mathbb{R}}^{d}\to\mathbb{T}^{r}({\mathbb{R}}^{d})$ is measurable and bounded on $\partial A$, then
@@
We apply (3) with the map $\tilde{f}:{\mathbb{R}}^{d}\to\mathbb{T}^{r+s}$, $\tilde{f}(z)=f(p_{A}(z))u_{A}(z)^{s}\mathds{1}\{z\in A^{\varepsilon}\setminus A\}$ (with the understanding that $\tilde{f}(z)=0$ if $z\in A$), which yields
@@
$\leq\frac{\omega_{d-i}}{d-i}\cdot|f|_{\partial A}\cdot\|\mu_{i}(A)\|(\operatorname{Nor}(A))\cdot\varepsilon^{d-i-1}\to 0\quad\text{ as }\varepsilon\to 0_{+},$ (3.4)
where $|f|_{\partial A}:=\sup\{|f(x)|:x\in\partial A\}<\infty$ and basic properties of the norm $|\cdot|$ on (symmetric) tensors were used, in particular (2.2). Moreover, by the dominated convergence theorem,
@@
since $r_{A}(x,u)>0$ for $(x,u)\in\operatorname{Nor}(A)$. Note that by [28, Proposition 4.1], $\mu_{d-1}(A;\cdot)$ is a nonnegative $\sigma$-finite Borel measure and
@@
Remark 3.8.
Relation (3.4) in the proof of Theorem 3.7 shows that
$\varepsilon^{-1}\left|\int_{0}^{\varepsilon}h_{i}(t)\,\mathrm{d}t\right|\leq C(d,A,f)\cdot\varepsilon,\quad i\in\{0,\ldots,d-2\},$
where $C(d,A,f)$ is a constant that depends only on the parameters in brackets. If $\operatorname{reach}(A)>0$, then $h_{d-1}(t)=h_{d-1}(0_{+})$ for $0<t<\operatorname{reach}(A)$, hence in this case the deviation
$\left|\frac{1}{\varepsilon}\int_{A^{\varepsilon}\setminus A}f(p_{A}(z))u_{A}(z)^{s}\,\mathcal{H}^{d}(\mathrm{d}z)-2\int_{\operatorname{Nor}(A)}f(x)u^{s}\,\mu_{d-1}(A;\mathrm{d}(x,u))\right|$ (3.6)
is at most of the order $\varepsilon$ as $\varepsilon\to 0_{+}$.
If $A=B^{2}(o,1)\cup B^{2}(2\Delta e,1)$, for some fixed $\Delta\in(0,1)$ and $e\in\mathbb{S}^{1}$, then (3.6) is still of the order $\varepsilon$. However, if $\Delta=1$, then (3.6) is of the order $\varepsilon^{1/2}$. Considering the epigraph of the function $x\mapsto\frac{1}{\beta}|x|^{\beta}$, $\beta>1$ fixed, and its reflection across the $e_{1}$-axis, one can construct a set $A_{\beta}$ for which the deviation (3.6) is of the order $\varepsilon^{1/\beta}$.
The limit in Theorem 3.7 involves only integrations over (parts of) the boundary of the compact set $A$, which corresponds to the contribution of the top order support measure $\mu_{d-1}(A;\cdot)$, which can be thought of as the surface area measure of $A$. However, in the proof we have to control the contributions of all support measures $\mu_{i}(A;\cdot)$, $i\in\{0,\ldots,d-2\}$. By essentially the same argument, we obtain the following variant of Theorem 3.7. The case $s=0$ of the next theorem is already covered by Theorem 3.7 with the domain of integration restricted to the complement of the set $A$.